# The Race between Technology and Woman:

### v8

@@ 1 Introduction
Over the decades, there have been substantial changes in the structure of wages and employment. The wage gap between male and female workers has narrowed in many countries, while that between skilled and unskilled workers has widened in some countries (Krueger, Perri, Pistaferri and Violante, 2010). At the same time, the share of labor in national income has declined in many countries (Karabarbounis and Neiman, 2014). During the period, there have been tremendous advances in information, communication, and computation technologies in industrialized nations. Such technological advances have been recognized as one of the causes for the widening of the skill wage gap and the decline in the labor share (Hornstein, Krusell and Violante, 2005; Acemoglu and Autor, 2011; Grossman and Oberfield, 2022; Karabarbounis, 2024).
@@
The direction and magnitude of changes in the relative wages and income shares of different types of labor due to the expansion of technologically advanced new equipment depend on the presence and degree of capital–skill or capital–gender complementarity. Male and female workers possess different sets of skills (Welch, 2000), as skilled and unskilled workers do (Griliches, 1969). Communication skills, in which women can possibly have a comparative advantage, are more likely to be complementary to information and communication technologies. Physical skills, in which men can possibly have a comparative advantage, are more likely to be substitutable with computation and automation technologies. The widespread use of new technologies might raise the demand for brains relative to brawn, which could result in a narrowing of the gender wage gap, a widening of the skill wage gap, and a decline in the income share of male unskilled labor.
@@
This study develops a unified framework to evaluate the effects of capital-embodied technological change on, as well as the contributions of factor inputs to, the relative wages and income shares of different types of labor. Our framework builds upon a multi-sector, multi-factor extension of the Greenwood, Hercowitz and Krusell (1997) model of investment-specific technological change. [^1] The terms, capital-embodied technological change and investment-specific technological change, are used interchangeably in this paper. We incorporate technological change embodied in new capital into an economy consisting of multiple sectors, where multiple consumption and investment goods are produced from multiple types of capital and labor. The presence and degree of capital–skill and capital–gender complementarities in the multi-sector economy are determined by the aggregate elasticities of substitution among different types of capital and labor. We show that the aggregate elasticities of substitution are sufficient statistics to measure the contributions of factor inputs to the relative wages and income shares of different types of labor. In this context, we generalize a model that explains changes in the wage premium to skill in terms of the race between technology and education (Tinbergen, 1974; Goldin and Katz, 2010) and a model that explains changes in the labor share in terms of capital deepening (Hicks, 1932; Elsby, Hobijn and Şahin, 2013). We further show that the aggregate elasticities of substitution are sufficient statistics to quantify the effects of capital-embodied technological change on the relative wages and income shares of different types of labor. In this context, we synthesize models that explain a rise in the skill premium or a decline in the labor share in terms of a fall in the relative price of capital equipment (Krusell, Ohanian, Rios-Rull and Violante, 2000; Karabarbounis and Neiman, 2014).
@@
We apply this framework to developed economies consisting of multiple sectors, where the production of goods and services requires multiple types of capital and labor. Given the fact that capital-embodied technological progress was accelerated by information processing equipment and software in the 1980s and 1990s (Cummins and Violante, 2002), we distinguish equipment related to information, communication, and computation technologies from other types of capital. We refer to the former as ICT capital and the latter as non-ICT capital. We also distinguish between male skilled, female skilled, male unskilled, and female unskilled labor to examine the sources of changes in gender and skill premia. Although it is notoriously difficult to identify the elasticities of substitution in the aggregate production function using time-series data (Diamond, McFadden and Rodriguez, 1978), it is possible to obtain the aggregate elasticities of substitution by estimating and aggregating the disaggregated production function parameters with disaggregated data (Sato, 1975; Oberfield and Raval, 2021).
@@
We estimate the sectoral production function parameters and obtain the aggregate elasticities of substitution among the two types of capital and the four types of labor using cross-country cross-industry data from OECD countries. The advantages of cross-country cross-industry data over firm-level data are that it enables us to measure equipment related to information, communication, and computation technologies and to classify labor by gender and skill. We find that ICT equipment is more complementary not only to skilled labor than unskilled labor but also to female labor than male labor. This finding is robust to the specification of the sectoral production function, the choice of instrumental variables, and the presence of labor market institutions, factor-biased technological change, and markups. We demonstrate that the expansion of ICT equipment contributes not only to the widening of the skill wage gap and the narrowing of the gender wage gap but also to a fall in the income share of male unskilled labor, which accounts for most of the decline in the labor share.
@@
We show that technological change embodied in ICT capital induces significant changes in the equilibrium relative wages and income shares of the four types of labor. On average in OECD countries, if all else were held constant, a fall in the relative price of ICT equipment between the years 1980 and 2005 would have reduced the skilled (unskilled) gender wage gap by 49 (10) percentage points and raised the male (female) skill wage gap by 27 (77) percentage points, while it would have raised the income share of male (female) skilled labor by 1.4 (2.0) percentage points and reduced the income share of male (female) unskilled labor by 7.1 (0.6) percentage points. However, whether the gender (skill) wage gap eventually narrows (widens) depends on the outcome of the race between progress in ICT and advances in female employment (educational attainment). We also show that the observed difference in the rate of decrease (increase) in the gender (skill) wage gap by skill (gender) can be explained in terms of the race between progress in ICT and advances in female employment (educational attainment).
@@ 2 Related Literature
This study adds to several strands of literature. First, it is related to the literature on technological change and wage inequality. This strand of literature highlights the importance of factor-biased technological change in accounting for changes in wage inequality in the United States and OECD countries, typically by estimating the aggregate production function. The widening of the skill wage gap is attributed to skill-biased technological change (Bound and Johnson, 1992; Katz and Murphy, 1992; Goldin and Katz, 2010), technological change embodied in capital equipment (Krusell et al., 2000; Taniguchi and Yamada, 2022), or sector-specific disembodied technological change (Buera, Kaboski, Rogerson and Vizcaino, 2022). The narrowing of the gender wage gap is attributed to gender-biased technological change (Johnson and Keane, 2013).
@@
Second, this study is related to the literature on the labor market impact of new technologies. This strand of literature indicates advances in information, communication, computation, and automation technologies as causes of changes in the structure of wages and employment in the United States and other OECD countries. The rate of increase in the wage bill share of skilled or female labor tends to be higher in ICT or computer-intensive industries (Autor, Katz and Krueger, 1998; Michaels, Natraj and Van Reenen, 2014; Raveh, 2015). The increased employment and wages of skilled or female labor are associated with a rise in occupations requiring nonroutine analytical and interactive tasks that are complementary to ICT and a decline in routine cognitive and manual tasks that are substitutable with ICT (Autor, Levy and Murnane, 2003; Black and Spitz-Oener, 2010). The widening of the skill wage gap is attributed to a relative decline in industries requiring routine tasks that are substitutable with automation technology (Acemoglu and Restrepo, 2022). The narrowing of the gender wage gap as well as the widening of the skill wage gap are attributed to the increased use of computers or the decreased price of capital equipment (Beaudry and Lewis, 2014; Burstein, Morales and Vogel, 2019; Caunedo, Jaume and Keller, 2023).
@@
Third, this study is related to the literature on technological change and the labor share of income. This strand of literature points to the relevance of technological change to the decline in the labor share in the United States and elsewhere. The decline in the labor share is attributed to capital-embodied technological change (Karabarbounis and Neiman, 2014; Eden and Gaggl, 2018), factor-biased technological change (Oberfield and Raval, 2021), or the adoption of robotic technology (Acemoglu and Restrepo, 2020).
This study contributes to these strands of literature in two ways. First, we provide the first estimates for the aggregate elasticities of substitution between ICT equipment and the four types of labor based on a multi-sector, multi-factor general equilibrium model, thereby measuring the contributions of specific factor inputs to changes in the relative wages and income shares of the four types of labor. Our results suggest that the expansion of ICT equipment underlies not only skill-biased but also gender-biased technological change noted in the literature. Second, we analytically derive and quantitatively evaluate the general equilibrium effects of technological change embodied in ICT equipment on the relative wages and income shares of different types of labor.
@@
Finally, from a methodological point of view, this study is related to Baqaee and Farhi (2019) and Oberfield and Raval (2021), who develop a framework to characterize and estimate the aggregate elasticity of substitution in an economy consisting of multiple industries. Relative to these studies, we consider a dynamic economy in which new capital accumulates more rapidly as technology progresses. Furthermore, we extend the framework to allow for endogenous labor supply and imperfect labor mobility across sectors. However, as in the wage inequality literature, we focus on the value-added output.
@@ 3 Theory
We develop a unified framework to evaluate the effects of capital-embodied technological change on, as well as the contributions of factor inputs to, the relative wages and income shares of different types of labor. Our framework builds upon the seminal work of Greenwood, Hercowitz and Krusell (1997) on investment-specific technological change. We consider a multi-sector economy in which multiple consumption and investment goods are produced in each sector from multiple types of capital and multiple types of labor, whereas Greenwood, Hercowitz and Krusell (1997) consider a two-sector economy in which consumption and investment goods are separately produced in different sectors from two types of capital and one type of labor. We first show how the relative wages and income shares of different types of labor can change according to factor inputs. We then show how they can change according to a more exogenous factor, technological change. Furthermore, we extend the framework to allow for endogenous labor supply and imperfect labor mobility across sectors. Our theoretical results require the specification of neither the aggregate nor sectoral production function.
3.1 Model
@@ 3.1.1 Household
where $p_{nt}$ is the price of goods produced in sector $n$, $w_{ft}$ is the wages of labor $f$, and $r_{ft}$ is the rental price of capital $f$. The supply of labor is assumed to be exogenously given, but this assumption will be relaxed in Section 3.5.
The household maximizes its utility by equalizing consumption across individuals: $c_{fnt}=c_{nt}$ and $c_{ft}=c_{t}$ for all $f$. The expenditure share of final goods $n$ in period $t$ is then defined as $\zeta_{nt}=p_{nt}y_{nt}/\sum_{n=1}^{N}p_{nt}y_{nt}=p_{nt}y_{nt}/p_{t}y_{t}$, where the aggregate price is $p_{t}=(\sum_{n=1}^{N}\theta_{n}^{1/(1-\eta)}p_{nt}^{-\eta/(1-\eta)})^{-(1-\eta)/\eta}$, and the aggregate output is $y_{t}=c_{t}+\sum_{f=1}^{F_{k}}(x_{ft}/q_{ft})$. The expenditure share can be derived as
@@
As in Greenwood, Hercowitz and Krusell (1997), investment-specific technology can be measured by the ratio of the aggregate price of investment goods, $p_{ft}$, to the aggregate price of consumption goods.
@@ 3.1.2 Firms
where $y_{nt}$ is the real output in sector $n$ in period $t$, $\ell_{fnt}$ is the quantity of labor $f\in\{1,\ldots,F_{\ell}\}$, $k_{fnt}$ is the quantity of capital $f\in\{1,\ldots,F_{k}\}$, $A_{nt}$ is sector-specific factor-neutral technology, and $\mathcal{F}_{n}(\cdot)$ is a sectoral production function. Below, to avoid notational clutter, we let $\ell_{fnt}$ denote not only labor for $f\in\{1,\ldots,F_{\ell}\}$ but also capital for $f\in\{F_{\ell+1},\ldots,F\}$ and $w_{ft}$ denote not only wages for $f\in\{1,\ldots,F_{\ell}\}$ but also the rental price of capital for $f\in\{F_{\ell+1},\ldots,F\}$.
The income share of factor $f$ in sector $n$ in period $t$ is defined as $\lambda_{\ell_{f}nt}=w_{ft}\ell_{fnt}/p_{nt}y_{nt}$. When the production function is homogeneous of degree one, the profits (8) can be written as $p_{nt}y_{nt}-y_{nt}\mathcal{\widetilde{C}}_{n}(w_{1t},\ldots,w_{Ft})/A_{nt}$, where $\mathcal{\mathcal{\widetilde{C}}}_{n}(\cdot)/A_{nt}$ is a unit cost function in sector $n$ in period $t$, and the demand for factor $f$ in sector $n$ in period $t$ can be written as $\ell_{fnt}=y_{nt}\widetilde{\mathcal{G}}_{fn}(w_{1t},\ldots,w_{Ft})/A_{nt}$, where $\widetilde{\mathcal{G}}_{fn}(\cdot)/A_{nt}$ is the demand for factor $f$ per unit output in sector $n$ in period $t$. By Shephard’s lemma, $\lambda_{\ell_{f}nt}=\partial\ln\mathcal{\mathcal{\widetilde{C}}}_{n}(w_{1t},\ldots,w_{Ft})/\partial\ln w_{ft}$ and
$\frac{\partial\ln\lambda_{\ell_{f}nt}}{\partial\ln w_{gt}}=\frac{\partial\ln\widetilde{\mathcal{G}}_{fn}\left(w_{1t},\ldots,w_{Ft}\right)}{\partial\ln w_{gt}}-\lambda_{\ell_{g}nt}+I_{\left(f,g\right)},$ (10)
where $I_{(f,g)}$ is the $fg$-th element of the $F\times F$ identity matrix. While $\lambda_{\ell_{f}nt}$ can be observed directly from data, $\partial\ln\lambda_{\ell_{f}nt}/\partial\ln w_{g}$ can be estimated after $\mathcal{F}_{n}(\cdot)$ is specified and $\widetilde{\mathcal{G}}_{fn}(\cdot)$ is derived.

### v9

@@ 1 Introduction
Over the decades, there have been substantial changes in the structure of wages and employment. The wage gap between male and female workers has narrowed in many countries, while that between skilled and unskilled workers has widened in some countries (Krueger, Perri, Pistaferri and Violante, 2010). At the same time, the share of labor in national income has declined in many countries (Karabarbounis and Neiman, 2014). During the period, there have been tremendous advances in information, communication, and computation technologies in industrialized nations. Such technological advances have been recognized as one of the causes for the widening of the skill wage gap and the decline in the labor share (Hornstein, Krusell and Violante, 2005; Acemoglu and Autor, 2011; Grossman and Oberfield, 2022; Karabarbounis, 2024). Nevertheless, changes in relative wages and income shares vary in direction, magnitude, or both across different types of labor classified by gender and skill, as described later.
@@
The direction and magnitude of changes in the relative wages and income shares of different types of labor due to the expansion of technologically advanced new equipment depend on the presence and degree of capital–skill or capital–gender complementarity. Male and female workers possess different sets of skills (Welch, 2000; Borghans, ter Weel and Weinberg, 2014; Cortes, Jaimovich and Siu, 2023; Rendall, 2024), as skilled and unskilled workers do (Griliches, 1969). Communication skills, in which women can possibly have a comparative advantage, are more likely to be complementary to information and communication technologies. Physical skills, in which men can possibly have a comparative advantage, are more likely to be substitutable with computation and automation technologies. The widespread use of new technologies might raise the demand for brains relative to brawn, which could result in not only a widening of the skill wage gap but also a narrowing of the gender wage gap and a decline in the income share of male unskilled labor.
@@
This study develops a unified framework to evaluate the effects of capital-embodied technological change on, as well as the contributions of factor inputs to, the relative wages and income shares of different types of labor. Our framework builds upon a multi-sector, multi-factor extension of the Greenwood, Hercowitz and Krusell (1997) model of investment-specific technological change. [^1] The terms, capital-embodied technological change and investment-specific technological change, are used interchangeably in this paper. We incorporate technological change embodied in new capital into an economy consisting of multiple sectors, where multiple consumption and investment goods are produced from multiple types of capital and labor. The presence and degrees of capital–skill and capital–gender complementarities in the multi-sector economy are determined by the aggregate elasticities of substitution among different types of capital and labor. We show that the aggregate elasticities of substitution are sufficient statistics for measuring the contributions of factor inputs to the relative wages and income shares of different types of labor. In this context, we generalize a model that explains changes in the wage premium to skill in terms of the race between technology and education (Tinbergen, 1974; Goldin and Katz, 2010) and a model that explains changes in the labor share in terms of capital deepening (Hicks, 1932; Elsby, Hobijn and Şahin, 2013). We derive analytical expressions that characterize not only the contribution of each input, holding the other inputs constant, but also the effects of capital-embodied technological change, allowing all inputs to vary. We further show that the aggregate elasticities of substitution are sufficient statistics for quantifying the effects of capital-embodied technological change on the relative wages and income shares of different types of labor. In this context, we synthesize models that explain a rise in the skill premium or a decline in the labor share in terms of a fall in the relative price of capital equipment (Krusell, Ohanian, Ríos-Rull and Violante, 2000; Karabarbounis and Neiman, 2014).
@@
We apply this framework to advanced economies consisting of multiple sectors, where the production of goods and services requires multiple types of capital and labor. Given the fact that capital-embodied technological progress was accelerated by information processing equipment and software in the 1980s and 1990s (Cummins and Violante, 2002), we distinguish equipment related to information, communication, and computation technologies from other types of capital. We refer to the former as ICCT capital and the latter as non-ICCT capital. We classify labor by gender and skill into four types: male skilled, female skilled, male unskilled, and female unskilled labor. In doing so, we analyze how technological change embodied in ICCT capital shapes the observed differences in the direction and magnitude of changes in the gender wage gap by skill, the skill wage gap by gender, and labor shares by gender and skill.
@@
Although it is notoriously difficult to identify the elasticities of substitution in the aggregate production function using time-series data (Diamond, McFadden and Rodriguez, 1978; Acemoglu, 2002), it is possible to obtain the aggregate elasticities of substitution by estimating and aggregating the disaggregated production function parameters with cross-sectional variation in disaggregated data (Sato, 1975; Oberfield and Raval, 2021). We estimate the sectoral production function parameters and obtain the aggregate elasticities of substitution between ICCT capital and the four types of labor using cross-country cross-industry data from OECD countries. The advantages of cross-country cross-industry data over firm-level data are that it enables us to measure equipment related to information, communication, and computation technologies and to classify labor by gender and skill. Our estimates of the aggregate elasticities of substitution indicate that ICCT capital is more complementary not only to skilled labor than unskilled labor but also to female labor than male labor. This finding is robust to the choice of instrumental variables and the presence of labor market institutions, factor-biased technological change, and price markups.
@@
We decompose changes in the relative wages and labor shares of the four types of labor into components attributable to the factor complementarity effect associated with progress in ICCT and the relative quantity effect associated with advances in educational attainment and female employment, and then further into components attributable to the two types of capital and the four types of labor. Our results suggest that the expansion of ICCT equipment contributes to the widening of the skill wage gap, the narrowing of the gender wage gap, and a fall in the income share of male unskilled labor that accounts for most of the decline in the labor share. The observed difference in the rate of decrease (increase) in the gender (skill) wage gap by skill (gender) can be explained in terms of the race between progress in ICCT and advances in female employment (educational attainment). From the 1980s to the 2000s in OECD countries, the rate of decline in the gender wage gap was greater among unskilled workers than among skilled workers, while the rate of increase in the skill wage gap was greater among male workers than among female workers. Our results suggest that these differences do not arise from an insufficient increase in the relative demand for female skilled labor, but rather from a substantial increase in the relative supply of female skilled labor. Furthermore, we quantify the general equilibrium effects of technological change embodied in ICCT equipment on the gender wage gap by skill, the skill wage gap by gender, and labor shares by gender and skill. On average across OECD countries, if all else were held constant, a fall in the relative price of ICCT equipment between the years 1980 and 2005 would have reduced the skilled (unskilled) gender wage gap by 49 (10) percentage points and raised the male (female) skill wage gap by 26 (77) percentage points, while it would have raised the income share of male (female) skilled labor by 1.4 (2.0) percentage points and reduced the income share of male (female) unskilled labor by 7.0 (0.7) percentage points.
@@ 2 Related Literature
This study adds to several strands of literature. First, it is related to the literature on technological change and wage inequality. This strand of literature highlights the importance of factor-biased technological change in accounting for changes in wage inequality in the United States and other OECD countries, typically by estimating the aggregate production function. The widening of the skill wage gap is attributed to skill-biased technological change (Bound and Johnson, 1992; Katz and Murphy, 1992; Goldin and Katz, 2010), technological change embodied in capital equipment (Krusell, Ohanian, Ríos-Rull and Violante, 2000; Maliar, Maliar and Tsener, 2022; Taniguchi and Yamada, 2022; Ohanian, Orak and Shen, 2023), or sector-specific disembodied technological change (Buera, Kaboski, Rogerson and Vizcaino, 2022). The narrowing of the gender wage gap is also attributed to gender-biased technological change (Johnson and Keane, 2013). Relative to these studies, we obtain the aggregate elasticities of substitution by estimating and aggregating sectoral production function parameters, and examine the sources of changes in the four wage gaps: skilled gender wage gap, unskilled gender wage gap, male skill wage gap, and female skill wage gap.
@@
Second, this study is related to the literature on the labor market impact of new technologies. This strand of literature indicates advances in information, communication, computation, and automation technologies as causes of changes in the structure of wages and employment in the United States and other OECD countries. The rate of increase in the wage bill share of skilled or female labor tends to be higher in ICCT or computer-intensive industries (Autor, Katz and Krueger, 1998; Michaels, Natraj and Van Reenen, 2014; Raveh, 2015). The increased employment and wages of skilled or female labor are associated with a rise in occupations requiring nonroutine analytical and interactive tasks that are complementary to ICCT and a decline in routine cognitive and manual tasks that are substitutable with ICCT (Autor, Levy and Murnane, 2003; Black and Spitz-Oener, 2010). The widening of the skill wage gap is attributed to a relative decline in industries requiring routine tasks that are substitutable with automation technology (Acemoglu and Restrepo, 2022). The widening of the skill wage gap is also attributed to the increased use of computers and other capital equipment, whereas the narrowing of the gender wage gap is not necessarily attributable to that increased use (Beaudry and Lewis, 2014; Burstein, Morales and Vogel, 2019; Caunedo, Jaume and Keller, 2023). [^2] Beaudry and Lewis (2014) consider an aggregate production function with a unit elasticity of substitution between computer equipment and skilled labor and an infinite elasticity of substitution between computer equipment and unskilled labor. Burstein, Morales and Vogel (2019) consider an aggregate production function, in which final output is produced from occupational outputs, which are in turn produced by production units, each using one type of capital (computing or other equipment) and one type of labor (classified by gender, education, and age). They specify an occupational production function with a unit elasticity of capital–labor substitution for all types of capital and labor. Caunedo, Jaume and Keller (2023) consider an aggregate production function, in which final output is produced from occupational outputs, which are in turn produced from capital and labor. They specify an occupational production function with an common elasticity of capital–labor substitution for all types of labor and an infinite elasticity of labor–labor substitution for all types of labor. Relative to these studies, we allow all four types of labor disaggregated by gender and skill to be imperfect substitutes for one another and substitutable to varying degrees with ICCT equipment within sectors, thereby permitting the separate identification of capital–skill and capital–gender complementarities. We impose neither specific values on the elasticities of substitution among the four types of labor nor on those between ICCT equipment and labor.
@@
Third, this study is related to the literature on technological change and the labor share of income. This strand of literature points to the relevance of technological change to the decline in the labor share in the United States and elsewhere. The decline in the labor share is attributed to capital-embodied technological change (Karabarbounis and Neiman, 2014; Eden and Gaggl, 2018), factor-biased technological change (Oberfield and Raval, 2021), or the adoption of robotic technology (Acemoglu and Restrepo, 2020). Relative to these studies, we examine the sources of changes in labor shares by gender and skill. We decompose changes in the income share of the four types of labor into components attributable to the relative quantity effect, whose sign depends on whether the aggregate elasticities of substitution exceed one, and to the factor complementarity effect, whose sign is determined by the relative magnitude of the aggregate elasticities of substitution.
@@
Finally, from a methodological point of view, this study is related to Baqaee and Farhi (2019) and Oberfield and Raval (2021), who develop a framework to characterize and estimate the aggregate elasticity of substitution in an economy consisting of multiple sectors. Relative to these studies, we consider a dynamic economy in which new capital accumulates more rapidly as technology progresses. We derive analytical expressions that decomposes changes in relative wages and income shares into components attributable to factor inputs and quantifies the general equilibrium effects of capital-embodied technological change on relative wages and income shares. We further extend the framework to allow for endogenous labor supply and imperfect labor mobility across sectors. As in the wage inequality literature, we focus on the value-added output.
This study contributes to these strands of literature in three ways. First, we provide the first estimates of the elasticities of substitution between ICCT equipment and four types of labor disaggregated by gender and skill. These estimates enable us to explain the observed differences in the direction and magnitude of changes in the gender wage gap by skill, the skill wage gap by gender, and labor shares by gender and skill in terms of technological change embodied in ICCT equipment. Second, we analytically derive and quantitatively evaluate the contributions of specific factor inputs to changes in the relative wages and income shares of different types of labor in OECD countries. Our results suggest that technological change embodied in ICCT equipment underlies gender-biased technological change, as well as skill-biased technological change noted in the literature. Finally, we analytically derive and quantitatively evaluate the general equilibrium effects of technological change embodied in ICCT equipment on the relative wages and income shares of different types of labor in OECD countries. Our results capture the effects of capital-embodied technological change more comprehensively because the elasticity of capital–labor substitution is allowed to vary within and across sectors.
@@ 3 Theory
We develop a unified framework to evaluate the effects of capital-embodied technological change on, as well as the contributions of factor inputs to, the relative wages and income shares of different types of labor. Our framework builds upon the seminal work of Greenwood, Hercowitz and Krusell (1997) on investment-specific technological change. We consider a multi-sector economy in which multiple consumption and investment goods are produced in each sector from multiple types of capital and multiple types of labor, whereas Greenwood, Hercowitz and Krusell (1997) consider a two-sector economy in which consumption and investment goods are separately produced in different sectors from two types of capital and one type of labor. We first show how the relative wages and income shares of different types of labor can change in response to each input, holding all other inputs constant. We then show how they can change in response to a fundamental determinant, technological change, allowing all inputs to vary. Furthermore, we extend the framework to allow for endogenous labor supply and imperfect labor mobility across sectors in Appendix A. Our theoretical results do not require specifying any aggregate or sectoral production function.
3.1 Environment
@@ 3.1.1 Household
where $p_{nt}$ is the price of goods produced in sector $n$, $w_{ft}$ is the wages of labor $f$, and $r_{ft}$ is the rental price of capital $f$. The supply of labor is assumed to be exogenously given, but this assumption will be relaxed in Appendix A.
The household maximizes its utility by equalizing consumption across individuals: $c_{fnt}=c_{nt}$ and $c_{ft}=c_{t}$ for all $f$. The expenditure share of final goods $n$ in period $t$ is then defined as $\zeta_{nt}=p_{nt}y_{nt}/\sum_{n=1}^{N}p_{nt}y_{nt}=p_{nt}y_{nt}/p_{t}y_{t}$, where the aggregate price of consumption goods is $p_{t}=(\sum_{n=1}^{N}\theta_{n}^{1/(1-\eta)}p_{nt}^{-\eta/(1-\eta)})^{-(1-\eta)/\eta}$, and the aggregate output is $y_{t}=c_{t}+\sum_{f=1}^{F_{k}}(x_{ft}/q_{ft})$. The expenditure share can be derived as
@@
As in Greenwood, Hercowitz and Krusell (1997), investment-specific technology can be measured by the ratio of the aggregate price of investment goods to the aggregate price of consumption goods.
@@
where the aggregate price of investment goods is $p_{ft}=(1/q_{ft})(\sum_{n=1}^{N}\theta_{n}^{1/(1-\eta)}p_{nt}^{-\eta/(1-\eta)})^{-(1-\eta)/\eta}$.
@@ 3.1.2 Firms
where $y_{nt}$ is the real output in sector $n$ in period $t$, $\ell_{fnt}$ is the quantity of labor $f\in\{1,\ldots,F_{\ell}\}$, $k_{fnt}$ is the quantity of capital $f\in\{1,\ldots,F_{k}\}$, $A_{nt}$ is sector-specific factor-neutral technology, and $\mathcal{F}_{n}(\cdot)$ is a sectoral production function.
Below, to avoid notational clutter, we let $\ell_{fnt}$ denote not only labor for $f\in\{1,\ldots,F_{\ell}\}$ but also capital for $f\in\{F_{\ell}+1,\ldots,F\}$ and $w_{ft}$ denote not only wages for $f\in\{1,\ldots,F_{\ell}\}$ but also the rental price of capital for $f\in\{F_{\ell}+1,\ldots,F\}$.
The income share of factor $f$ in sector $n$ in period $t$ is defined as $\lambda_{\ell_{f}nt}=w_{ft}\ell_{fnt}/p_{nt}y_{nt}$. When the production function is homogeneous of degree one, the profits (8) can be written as $p_{nt}y_{nt}-y_{nt}\mathcal{\widetilde{C}}_{n}(w_{1t},\ldots,w_{Ft})/A_{nt}$, where $\mathcal{\mathcal{\widetilde{C}}}_{n}(\cdot)/A_{nt}$ is a unit cost function in sector $n$ in period $t$, and the demand for factor $f$ in sector $n$ in period $t$ can be written as $\ell_{fnt}=y_{nt}\widetilde{\mathcal{G}}_{fn}(w_{1t},\ldots,w_{Ft})/A_{nt}$, where $\widetilde{\mathcal{G}}_{fn}(\cdot)/A_{nt}$ is the demand for factor $f$ per unit output in sector $n$ in period $t$. By Shephard’s lemma, $\lambda_{\ell_{f}nt}=\partial\ln\mathcal{\mathcal{\widetilde{C}}}_{n}(w_{1t},\ldots,w_{Ft})/\partial\ln w_{ft}$ and $\partial\ln\lambda_{\ell_{f}nt}/\partial\ln w_{gt}=\partial\ln\widetilde{\mathcal{G}}_{fn}\left(w_{1t},\ldots,w_{Ft}\right)/\partial\ln w_{gt}-\lambda_{\ell_{g}nt}+I_{\left(f,g\right)},$where $I_{(f,g)}$ is the $fg$-th element of the $F\times F$ identity matrix. While $\lambda_{\ell_{f}nt}$ can be observed directly from data, $\partial\ln\lambda_{\ell_{f}nt}/\partial\ln w_{g}$ can be estimated after $\mathcal{F}_{n}(\cdot)$ is specified and $\widetilde{\mathcal{G}}_{fn}(\cdot)$ is derived.